A 5.0kg mass has its velocity change 8 m/s to 2 m/s east. Find thee objects change in momentum

1
Expert's answer
2022-05-03T14:21:23-0400
"\Delta p=m\Delta v\\
=\rm 5.0\: kg\times (8\: m/s-2\: m/s)=30\: kg\cdot m/s"